# Mobile Edge Generation-Enabled Digital Twin: Architecture Design and Research Opportunities

### v1

@@ Abstract
A novel paradigm of mobile edge generation (MEG)-enabled digital twin (DT) is proposed, which enables distributed on-device generation at mobile edge networks for real-time DT applications. First, an MEG-DT architecture is put forward to decentralize generative artificial intelligence (GAI) models onto edge servers (ESs) and user equipments (UEs), which has the advantages of low latency, privacy preservation, and individual-level customization. Then, various single-user and multi-user generation mechanisms are conceived for MEG-DT, which strike trade-offs between generation latency, hardware costs, and device coordination. Furthermore, to perform efficient distributed generation, two operating protocols are explored for transmitting interpretable and latent features between ESs and UEs, namely sketch-based generation and seed-based generation, respectively. Based on the proposed protocols, the convergence between MEG and DT are highlighted. Considering the seed-based image generation scenario, numerical case studies are provided to reveal the superiority of MEG-DT over centralized generation. Finally, promising applications and research opportunities are identified.
@@ I Introduction
With recent decade’s booming developments in artificial intelligence (AI) and wireless networks, digital twin (DT) has emerged as a promising approach for creating virtual replications of real-world physical objects and processes [1]. By revolutionizing the fashions of modelling, monitoring, analytics, and decision-making for intricate real-world systems, DT is anticipated to transform diverse facets of human society, spanning industry, education, agriculture, healthcare, and beyond. To maintain immersive and real-time evolving virtual duplicates for physical entities, DT requires ubiquitous connectivity, high-speed transmission, low latency, and intensive computing on mobile devices [2, 3]. To meet these challenges, a prospective solution is to integrate DT with the forthcoming sixth generation (6G) wireless networks, which leads to a win-win option. On the one hand, 6G wireless networks enable real-time synchronized DTs by incorporating new communication technologies (e.g., millimetre-wave/terahertz communications and satellite communications), enhanced infrastructures (e.g., integrated sensing and communications), as well as cloud/edge intelligence [3]. On the other hand, DT provides powerful tools to realize precise modelling (e.g., channel modelling) and self-sustaining management of 6G wireless networks [4].
@@
While holding immense potentials, establishing and maintaining high-fidelity virtual twins suffer from serious challenges and significant costs, particularly for real-time mobile DT applications. Fortunately, benefiting from advances in generative AI (GAI) and large-scale Transformer models, AI generated contents (AIGC) has unlocked new pathways to generate high-fidelity and vivid digital contents in virtual worlds [5]. The great success of GAI models, such as Generative Pre-training (GPT) [6] and Stable Diffusion [7], has spawned numerous AIGC applications in Natural Language Processing (NLP) and computer vision, involving text generation, text-to-image generation, image translation/synthesis, and video generation. With the impressive capabilities to comprehend data patterns and mimic human creations, AIGC can automatically and seamlessly map real-world contents into virtual twins. Therefore, AIGC has been regarded as a key enabler of DT for smart content transformation and enhancement [5]:
  • Content Transformation: Constructing virtual twins of physical worlds requires seamless transformations across multi-modal data (e.g., multilingual text, audio, image, video, and 3d point clouds) via AIGC.
  • Content Enhancement: The completeness, quality, and customization of digital contents should be improved by AIGC. For instance, AIGC can compensate for missing information, augment image/video quality, mitigate distortion/noise, edit content via human-computer interaction, and so on.
Despite these promising benefits, the size of GAI models has increased to 175 billions of parameters in GPT-3 from several hundreds of millions of parameters in GPT [6]. Following the stunning generative performance presented by GPT-3-style models, the development of large models has become a popular pursuit in AIGC domain. In the most recent GPT-4 model, the number of parameters has grown to an unprecedented 1.7 trillion, which is nearly 1000 times larger than GPT-3. However, training and executing a large-scale GAI model rely on high computing powers from massive GPUs [8]. Therefore, the deployments of large-scale GAI models on mobile devices suffer from overwhelming hardware complexity and serious energy concerns, which limits their applicabilities in real-time DT. To relieve UEs’ computational burdens, mobile edge networks can be deployed for DT, where the generation tasks can be offloaded from hardware-constrained user equipments (UEs) to high-performance edge servers (ESs). However, as massive mobile devices may require AIGC capabilities in real-time DT applications, transmitting huge-volume data for centralized generation is time-consuming and privacy-sensitive. This requires edge learning and inference design for large-scale GAI models in DT systems.
@@
Recently, the concept of mobile edge generation (MEG) has been proposed [9]. Specifically, MEG decentralizes GAI models into sub-models deployed on UEs and ESs. By doing so, distributed generation can be achieved by transmitting intermediate features of GAI sub-models between the heterogeneous hardware devices, without exchanging the raw data. In comparison to centralized generation, MEG can significantly reduce the latency and hardware costs for mobile UEs to access large-scale GAI models. Therefore, it enables an innovative paradigm to exploit large-scale GAI models like GPT-4 and DALL-E in real-time DT systems. However, to the best of our knowledge, the integration of MEG and DT has not been explored yet. To fill in this gap, we propose a novel MEG-enabled DT (MEG-DT) architecture in this article, targeted at realizing efficient on-device generation at the mobile edge networks of DT. Our main contributions can be summarized as follows.
@@
II MEG-DT: Architecture, Key Features, and Mechanisms
@@
In this section, we first propose a novel MEG-DT architecture. Then, we highlight the key advantages of MEG-DT and present tailored single-user and multi-user generation mechanisms.
II-A The Proposed MEG-DT Architecture
As shown in Fig. 1, we consider a mobile edge network for MEG-DT that consists of heterogeneous hardware devices, including high-performance ESs and low-cost UEs. To realize on-device mobile generation, large-scale GAI models are decentralized into collaborative GAI sub-models that are jointly deployed on ESs and UEs. By transmitting the features extracted from GAI sub-models between heterogeneous hardware devices, high-quality generation can be realized by jointly exploiting datasets, computational resources, and customized sub-models on both sides. In this way, MEG can bridge between ESs and mobile UEs as an immediate layer for DT construction and applications. With the tailored deployment and design of the distributed GAI sub-models, the proposed MEG-DT architecture can realize real-time content translation and enhancement to provide on-device DT services, such as multi-modal sensing, virtual replica construction, data analytics, performance prediction, and decision-making.
II-B Key Advantages of MEG-DT
The key advantages of MEG can be summarized as follows.
  • Low-Latancy Feature Compression and Transmission: Unlike centralized generation that directly exchanges raw data between UEs and ESs, MEG transmits features constructed from sub-models, which encode and compress the raw data into low-dimensional tensors. Therefore, the communication overheads and end-to-end generation latency required by MEG can be significantly reduced.
  • Distributed Model Deployment and Generation: MEG enables the distributed deployment of a large-scale GAI model on mobile UEs and the high-performance ES. With this innovative paradigm, the computing resources and local datasets from heterogeneous hardware devices can be efficiently orchestrated in various application scenarios.
  • Privacy Preservation and Security: Since UEs/ES can selectively expose and share their private data with each other, MEG can preserve the data privacy for various DT applications. Furthermore, by transmitting the implicit features during content generation, the system security can be significantly improved.
  • Individual-Level Customized Generation: Relying on the MEG paradigm, the model parameters, hyperparameters, and neural layer structures of GAI sub-models deployed on different UEs can be customized according to their individual preferences, particular generation purposes, and specific hardware constraints.
@@ II-C Single-User and Multi-User Generation for MEG-DT
Based on the proposed MEG-DT architecture, we discuss the single-user and multi-user generation mechanisms.
@@ II-C1 Single-User Generation
As shown in Fig. 2, the proposed MEG-DT architecture mainly contains four single-user generation mechanisms. Specifically, UE-to-ES (U2E) generation and ES-to-UE (E2U) generation enable feature encoding and content generation at ES and UE, respectively. Furthermore, sequential ES-and-UE (SEU) generation and parallel ES-and-UE (PEU) generation enable feature encoding and content generation on both sides, as discussed as follows.
  • U2E Generation: By sensing physical environment, UE encodes the collected data via its GAI sub-model to construct low-dimensional features. After receiving features transmitted from UE, ES generates digital contents via its GAI sub-model. A typical use case of U2E generation in DT is to generate/refresh the virtual replica at ES using the multi-modal sensing data (e.g., text, voice, camera, radar/LiDAR, inertial measurement unit (IMU)) from UE.
  • E2U Generation: ES encodes the local data and/or the inference results via its GAI sub-model and transmits the encoded features to target UE. By decoding the received features, UE further reconstructs inference results and generates desired contents. For example, UE can generate desired predictions by retrieving data (e.g., high-precision 3D map and expert knowledge) from the server.
  • SEU Generation: SEU generation allows ES and UE to sequentially transmit encoded features and generate contents, enabling virtual-reality synchronization. UE first transmits extracted features to ES in the uplink. Next, ES generates contents and updates server dataset, and then transmits downlink features to assist in the generation on UE. This mechanism can be employed when UE has very limited computing/storage capacities and requires ES to handle the vast majority of computing tasks.
  • PEU Generation: PEU generation allows UE and ES to simultaneously perform feature encoding, transmission, and content generation in a parallel manner, as shown in Fig. 2(d). Specifically, a feature gate is learned to integrate features from both sides. Local copies of the feature gate are deployed on both ES and UE, which create representations of a synchronized and shared view for parallel generation. Compared to SEU generation, this mechanism can reduce end-to-end generation latency for delay-sensitive DT applications, but requires higher hardware costs for UEs.
@@ II-C2 Multi-User Generation
Since each UE can only perceive their surrounding environment, the application of MEG-DT may require interactions between multiple UEs. By integrating UEs’ local features, a global view of physical/digital world can be established for immersive generation and cooperative decision-making. To reap the benefits, we conceive three multi-user generation mechanisms for MEG-DT, which can realize distinct levels of interactions and coordination between UEs, as shown in Fig. 3. Here, we focus on multi-user interactions for UE-side generation. The BS-side interaction can be achieved similarly.
@@
  • Coordinated Generation via Fused Features: The third level enables UEs to thoroughly fuse their local features at the nearby ES for enhanced coordination. Exploiting the fused features, fully synchronized and coordinated multi-user generations can be realized. This mechanism can be employed when UEs require a comprehensive picture of physical world, while tolerating increased latency and computing complexity.
@@ III MEG-DT: Principle and Interplays
To realize efficient distributed generation during the single-user and multi-user generations, a key issue is to effectively extract low-dimensional features that encode local data at specific ESs/UEs. In this section, we will introduce the main principles to extract useful features by exploring two protocols, i.e., sketch-based generation and seed-based generation. Then, the convergence between MEG and DT will be discussed.
@@ III-A1 Sketch-Based Generation
Sketch typically represents a simplified image or drawing that only retains object outlines and essential spatial characteristics of the original image. As it provides a low-complexity visual representation of image, it has been widely applied in image retrievals and analytics. Derived from this concept in computer vision, in the considered MEG-DT settings, we refer to the sketch as a low-complexity and interpretable abstraction of the raw data (such as text, image, and 3D maps), which can be used for a wide rage of generation tasks.
To efficiently perform sketch-based generation, ESs and UEs can jointly train domain-specific encoder and decoder that bridge between the distributed GAI sub-models, as shown in Fig. 4(a). Specifically, the encoder learns to extract a sketch of the raw data, thus preserving important information while reducing data volumes. Then, by transmitting the obtained sketch to the decoder deployed on another hardware equipment, the details will be reconstructed for high-quality generations. To efficiently compress and reconstruct data, the encoder and decoder can inherit the architectures from current GAI models for image-to-image translation [10], e.g., generative adversarial networks (GANs), variational auto-encoder (VAE), and diffusion models. These models can be further tailored for different generation purposes. During generation, text prompts can be combined to control the content editing and synthesis, thus flexibly synthesizing virtual/realistic datasets on ES and UEs. In general, sketch-based generation is easy to implement, guide, and examine.
Figure 4: Diagram of sketch-based generation and seed-based generation protocols.
@@ III-A2 Seed-Based Generation
Instead of transmitting sketches that have explicit interpretations, another promising approach is to transform the high-dimensional data into latent features without intuitive explanations. These latent features are known as seeds. In comparison to the sketch-based generation that creates contents on the original feature space of the physical-world data, seed-based generation can deal with a more compact latent feature space. By identifying the underlying dependence of data implicitly, seed-based generation is more efficient to process large-scale or sequential data, such as time-varying sensing observations and videos.
A simple implementation of seed-based generation is to split an existing GAI model into two sub-model. Thus, seeds can be constructed by the hidden layer that splits the original GAI model. During forward/backward propagation, only features/gradients of the split hidden layer need to be transmitted between different devices. In practice, pretrained GAI models (e.g., Stable Diffusion) can be directly split into encoder and decoder for seed-based generation without additional training process. Nevertheless, the achievable end-to-end generation quality is still limited by practical compression performance and noisy wireless channels. To fully exploit the potentials of seed-based generation and overcome these limitations, advanced MEG-DT strategies should be developed to systematically realize efficient seed construction, compression, and source-channel coding, thus enhancing on-device generation quality with lower system complexity and overheads, as discussed as follows.
@@
Figure 5: The suitable MEG protocols for different generation tasks.

### v3

@@ Abstract
A novel paradigm of mobile edge generation (MEG)-enabled digital twin (DT) is proposed, which enables distributed on-device generation at mobile edge networks for real-time DT applications. First, an MEG-DT architecture is put forward to decentralize generative artificial intelligence (GAI) models onto edge servers (ESs) and user equipments (UEs), which has the advantages of low latency, privacy preservation, and individual-level customization. Then, various single-user and multi-user generation mechanisms are conceived for MEG-DT, which strike trade-offs between generation latency, hardware costs, and device coordination. Furthermore, to perform efficient distributed generation, two operating protocols are explored for transmitting interpretable and latent features between ESs and UEs, namely sketch-based generation and seed-based generation, respectively. Based on the proposed protocols, the convergence between MEG and DT are highlighted. Considering the seed-based image generation scenario, numerical case studies are provided to reveal the superiority of MEG-DT over centralized generation. Finally, promising applications and research opportunities are identified. Code is available at https://github.com/xiaoxiaxusummer/MEG_DT.
@@ I Introduction
With recent decade’s booming developments in artificial intelligence (AI) and wireless networks, digital twin (DT) has emerged as a promising approach for creating virtual replications of real-world physical objects and processes [1]. By revolutionizing the fashions of modelling, monitoring, analytics, and decision-making for intricate real-world systems, DT is anticipated to transform diverse facets of human society, spanning industry, education, agriculture, healthcare, and beyond. To maintain real-time evolving duplications of physical entities, DT requires ubiquitous connectivity, high-speed transmission, low latency, and intensive computing on mobile devices [2, 3]. To meet these challenges, a prospective solution is to integrate DT with the forthcoming sixth generation (6G) wireless networks, which leads to a win-win option. On the one hand, 6G wireless networks enable real-time synchronized DTs by incorporating new communication technologies (e.g., millimetre-wave/terahertz communications and satellite communications), enhanced infrastructures (e.g., integrated sensing and communications), as well as cloud/edge intelligence [3]. On the other hand, DT provides powerful tools to realize precise modelling (e.g., channel modelling) and self-sustaining management of 6G wireless networks [4].
@@
While holding immense potentials, establishing and maintaining high-fidelity virtual twins suffer from significant manpower and economic costs, particularly for real-time DT constructions. Fortunately, benefiting from advances in generative AI (GAI) and large-scale Transformer models, AI generated contents (AIGC) has unlocked new pathways to generate high-fidelity and vivid digital contents in virtual worlds [5]. The great success of GAI models, such as Generative Pre-training (GPT) [6] and Stable Diffusion [7], has spawned numerous popular AIGC applications in information industries, involving text generation, text-to-image generation, image translation/synthesis, and video generation. With the impressive capabilities to comprehend data patterns and to mimic artistic creations, AIGC can automatically and seamlessly map real-world contents into virtual twins, thus enabling smart content transformation and enhancement for DT [5]. However, the stunning generative performance relies on large-scale GAI models [7, 6]. Specifically, the popular GPT-3 model requires 175 billions parameters [6]. Moreover, the number of parameters in the most recent GPT-4 model has grown to an unprecedented 1.7 trillion, which is nearly 1000 times larger than GPT-3. Since deploying these large-scale GAI models on resource-limited mobile devices leads to an unaffordable hardware complexity [8], their applications in DT systems require powerful cloud services from distant datacenter.
@@
To reduce the access latency of remote cloud servers and achieve ubiquitous AIGC services, mobile edge networks can be deployed for real-time DT systems, which offload generation tasks from resource-limited user equipments (UEs) to nearby edge servers (ESs). However, as DT systems comprise massive devices with intensive generation requests, it is time-consuming and privacy-sensitive to transmit huge-volume data for accessing centralized GAI models at ESs. Against this background, the concept of mobile edge generation (MEG) has been recently proposed [9]. Specifically, MEG decentralizes GAI models into sub-models deployed on UEs and ESs,which provides a tailored edge AI solution for mobile AIGC. As a result, distributed generation can be achieved by transmitting intermediate features of GAI sub-models between different hardware devices without exchanging the raw data. In comparison to centralized generation, MEG can significantly reduce the latency and hardware costs for mobile UEs to access large-scale GAI models. Therefore, it enables an innovative paradigm to exploit large-scale GAI models like GPT-4 and DALL-E in real-time DT systems. However, to the best of our knowledge, the integration of MEG and DT has not been explored yet. To fill in this gap, we propose a novel MEG-enabled DT (MEG-DT) architecture in this article, targeted at realizing efficient generation on edge devices for DT systems. Our main contributions can be summarized as follows.
@@
II MEG-DT: Architecture, Key Advantages, and Mechanisms
In this section, we first propose a novel MEG-DT architecture. Then, we highlight key advantages of MEG-DT and present single-user and multi-user generation mechanisms.
II-A The Proposed MEG-DT Architecture
As shown in Fig. 1, we consider a mobile edge network for MEG-DT that consists of heterogeneous hardware devices, including high-performance ESs and low-cost UEs. MEG provides mobile AIGC services across edge networks to construct and maintain DT. The key idea of MEG is to decompose the large-scale GAI models into decentralized GAI sub-models deployed on ESs and UEs, respectively. Instead of transmitting raw data, ESs and UEs only need to exchange low-dimension features extracted by GAI sub-models for content generation. Hence, distributed datasets, computing resources, and radio resources across mobile edge networks can be efficiently used. Relying on the distributed generation, MEG-DT enables autonomous and cost-effective content transformation/editing, content enhancement, and content synthesis for virtual twins’ creation and maintenance, as specified as follows.
  • Content Transformation and Editing: MEG-DT can transform the physical-world data collected by mobile UEs’ sensors (e.g., voices, photos/videos, radar/LiDAR point clouds, and inertial measurement unit (IMU) records) into virtual replicas. Thus, the digital mapping to real-world entities can be created automatically, e.g., for smart city DT creations. The created virtual contents can be further edited by specific prompts according to users’ intentions.
  • Content Enhancement: MEG-DT can improve the completeness and quality of virtual contents, such as compensating for missing information, augmenting image/video quality, and mitigating distortion. This can significantly enhance virtual replicas’ fidelity and user experiences for immersive DT applications (e.g., cloud game and extended reality (XR)).
  • Content Synthesis: MEG-DT can synthesize multi-modal virtual and/or reality data from different edge devices. This enables the virtual-reality synchronization and the real-time interactions between virtual/reality entities for interactive DT applications (e.g., interconnected robotics and Metaverse).
@@ II-A The Proposed MEG-DT Architecture
II-B Key Advantages
Key advantages of MEG-DT can be summarized as follows.
  • Low-Latency Feature Transmission: MEG transmits features constructed from GAI sub-models, which encode and compress the raw data into low-dimension representations. Compared to centralized generation that directly exchanges raw data between UEs and ESs, the communication overheads and end-to-end generation latency required by MEG-DT can be significantly reduced.
  • Efficient Edge Resource Utilization: MEG enables distributed GAI model deployment and flexible collaborations between multiple UEs and ESs. Therefore, the computing/radio resources and local datasets from heterogeneous devices can be efficiently orchestrated to enhance service efficiency in various application scenarios.
  • Privacy and Security Preservation: Unlike centralized generation that directly transmits users’ data, MEG-DT allows UEs/ESs to transmit implicit features that encodes users’ private data to protect against eavesdropping. Therefore, MEG-DT can preserve the data privacy and improve system security for various DT applications.
  • Individual-Level Customized Generation: Relying on MEG-DT, the neural parameters and structures of UEs’ local sub-models can be customized according to their individual preferences, personal generation purposes/demands, and particular hardware constraints.
@@ II-C Single-User and Multi-User Generation for MEG-DT
Based on the proposed MEG-DT architecture, we discuss single-user and multi-user generation mechanisms for different scenarios, which strike trade-offs between the latency, hardware complexity, and device coordination.
@@ II-C1 Single-User Generation
As shown in Fig. 2, the proposed MEG-DT architecture mainly contains four single-user generation mechanisms. Specifically, ES-to-UE (E2U) generation and UE-to-ES (U2E) generation performs feature encoding at the UE and the ES, respectively. Furthermore, sequential ES-and-UE (SEU) generation and parallel ES-and-UE (PEU) generation enable feature encoding and content generation on both sides, as discussed as follows.
  • E2U Generation: The ES first receives uplink (UL) task requests from the UE and performs inference. Then, low-dimension features are extracted from inference results and transmitted to the UE via downlink (DL) channels. Using the received features, the UE further decodes desired contents. This enables the UE to perform prompt-based generations by exploiting large-scale GAI sub-model and databases deployed at the ES.
  • U2E Generation: The UE encodes the collected sensing data into low-dimension features. By receiving features from UL transmissions, the ES generates digital contents. This allows the ES to create and update virtual replicas exploiting the multi-modal sensing data from UE.
  • SEU Generation: SEU generation allows the ES and the UE to sequentially perform feature encoding and transmissions for content editing/synthesis. The UE first encodes local data and uploads the encoded features to the ES by UL transmissions. Relying on this, the ES creates contents and transmits extracted features via DL channels to facilitate UE-side generation. This mechanism can be employed to edit or synthesize contents when the UE has very limited computing and storage capacities.
  • PEU Generation: PEU generation enables the UE and the ES to simultaneously transmit features in both UL and DL channels, and thus generate contents in parallel, as shown in Fig. 2(d). A feature gate is learned to integrate features from both sides. Local copies of the feature gate are deployed at both the ES and the UE, thus enabling a shared view for synchronized and parallel generation. Compared to SEU generation, PEU enables fast data synthesis and ES-UE interactions for delay-sensitive DT applications, but requires higher hardware costs for UEs.
@@ II-C2 Multi-User Generation
Maintaining real-time DTs requires devices’ interaction/cooperation. By integrating UEs’ local features extracted from the surrounding environment, a global view of physical/digital world can be established for immersive experiences and cooperative decision-making. However, devices’ interactions will introduce extra transmission latency. To strike a balance, we conceive three multi-user generation mechanisms, which can realize distinct levels of interaction/coordination and latency for MEG-DT, as shown in Fig. 3. We focus on UE-side cooperation and generation in this part. The ES-side cooperation can be achieved similarly.
@@
  • Coordinated Generation via Fused Features: The third level enables UEs to thoroughly fuse their local features at the nearby ES for enhanced coordination. Exploiting the fused features, fully synchronized and coordinated multi-user generations can be realized. This mechanism can be employed when UEs require large-scale data synthesis and mutual interactions, while tolerating increased latency and computing complexity.
@@ III MEG-DT: Principle and Interplays
To realize efficient distributed generation during the single-user and multi-user generations, a key issue is to effectively extract low-dimension features that encode local data of ESs/UEs. In this section, we will introduce the main principles to extract useful features by exploring two protocols, i.e., sketch-based generation and seed-based generation. Then, the convergence between MEG and DT will be discussed.
@@ III-A1 Sketch-Based Generation
Sketch typically refers to a simplified drawing that only retains object outlines and essential spatial characteristics of the original image. Derived from this concept in computer vision, we refer to sketch as a low-complexity but human-understandable representation for various types of data in DT. To name a few, sketches can represent summaries of texts, contours of 3D maps, snapshots of videos, and so on. Sketch-based generation compresses data into sketches for distributed generation. As such, users can gain a fast understanding of the generated contents before recovering the high-dimension data, thus easily examining and guiding creations during distributed generation.
To efficiently perform sketch-based generation, ESs and UEs can jointly train domain-specific sketch encoder and decoder to bridge between the distributed GAI sub-models, as shown in Fig. 4(a). Specifically, the encoder learns to extract a sketch of the raw data, thus preserving important information while reducing data volumes. Then, by transmitting the obtained sketch to the decoder deployed on another hardware equipment, the details will be reconstructed for high-quality generations. To efficiently compress and reconstruct data, the encoder and decoder can inherit the architectures from current GAI models for image-to-image transformation [10], e.g., generative adversarial networks (GANs), variational auto-encoder (VAE), and diffusion models. These models can be further tailored for different generation purposes. During generation, text prompts can be combined to control the content editing and synthesis, thus flexibly synthesizing virtual/realistic datasets on ES and UEs.
Figure 4: Sketch-based and seed-based generation protocols.
@@ III-A2 Seed-Based Generation
Instead of transmitting sketches that have explicit interpretations, another promising approach is to transform the high-dimension data into latent features without intuitive explanations. These latent features are known as seeds. In comparison to the sketch-based generation that creates contents on the original feature space of the physical-world data, seed-based generation can deal with a more compact latent feature space. By identifying the underlying dependence of data implicitly, seed-based generation is more efficient to process large-scale or sequential data, such as time-varying sensing observations and videos.
A simple implementation of seed-based generation is to directly split an existing GAI model (e.g., Stable Diffusion) into two sub-models, and then construct seeds by the intermediate hidden layers. Similar to existing edge AI solutions, the edge-device co-inference can be achieved. However, different from classification/regression tasks, MEG targets at content generation, where the end-to-end generation quality is sensitive to wireless channel noises and feature compression. To fully exploit the potentials of seed-based generation and overcome these limitations, advanced MEG-DT strategies should be developed to realize efficient seed construction, compression, and source-channel coding, thus enhancing on-device generation quality while relieving system complexity and overheads, as discussed as follows.
@@
Figure 5: Suitable MEG mechanisms/protocols for various use cases and application scenarios.